OpenStudy (anonymous):

divide 5-2i/3+6i. Please help!

OpenStudy (anonymous):

multiply top and bottom by the conjugate of the denominator

OpenStudy (anonymous):

the conjugate of \(a+bi\) is \(a-bi\) and this works because \((a+bi)(a-bi)=a^2+b^2\) a real number

OpenStudy (anonymous):

in your case it will be \[\frac{5-2i}{3+6i}\times \frac{3-6i}{3-6i}=\frac{(5-2i)(3-6i)}{9+36}\]

OpenStudy (anonymous):

and you finish it up from there?

OpenStudy (anonymous):

i got 15-36i+12i^2/9-6i

OpenStudy (anonymous):

the denominator should have no \(i\) in it that is what you are trying to achieve

OpenStudy (anonymous):

i^2 is 1 right?

OpenStudy (anonymous):

\[(3+6i)(3-6i)=3^2+6^2=9+36=45\]

OpenStudy (anonymous):

no \(i=\sqrt{-1}\) so \(i^2=-1\) not \(1\)

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

make sure it is clear what you are trying to do in this problem your goal is to divide, which really means turn \[\frac{5-2i}{3+6i}\] into standard form as \(a+bi\)

OpenStudy (anonymous):

i wrote the steps above, although you still have to compute the numerator

OpenStudy (anonymous):

i got 1/15 - 36square root of -1/45

OpenStudy (anonymous):

there is not square root in it

OpenStudy (anonymous):

only i instead of square root

OpenStudy (anonymous):

first part is right, it is \(\frac{1}{15}\) for the real part

OpenStudy (anonymous):

36i over 45?

OpenStudy (anonymous):

oh i guess the whole thing you wrote is correct

OpenStudy (anonymous):

except you should write \[\frac{1}{15}-\frac{4}{5}i\]

OpenStudy (anonymous):

thank you so much! you are my hero now!

OpenStudy (anonymous):

my pleasure